# Align Your Intents: Offline Imitation Learning via Optimal Transport

### v1

@@ 5.2 Baselines
  • IQL Kostrikov et al. (2021) is state-of-the-art offline RL algorithm, which avoids querying out of the distribution actions by viewing value function as a random variable, where upper bound of uncertanity is controlled through expectile of distribution. In our experiments, evaluation is made using ground-truth reward from D4RL tasks.
  • OTR Luo et al. (2023) is a reward function algorithm, where reward signal is based on optimal transport distance between states of expert demonstration and reward unlabeled dataset.
  • CLUE Liu et al. (2023) learns VAE calibrated latent space of both expert and agent state-action transitions, where intrinsic rewards can be defined as distance between agent and averaged expert transition representations.
  • IQ-Learn Garg et al. (2021) is an imitation learning algorithm, which implicitly encodes into learned inverse Q-function rewards and policy from expert data.
  • SQIL Reddy et al. (2019) proposes to learn soft Q-function by setting expert transitions to one and for non-expert transitions to zero.
  • ORIL Zolna et al. (2020) utilizes discriminator network which distinguishes between optimal and suboptimal data in mixed dataset to provide reward relabelling through learned discriminator.
  • SMODICE Ma et al. (2022a) offline state occupancy matching algorithm, which solves the problem of IL from observations through state divergence minimization by utilizing dual formulation of value function.
5.3 Results
In this section, we show the results for the following offline RL settings: (1) the imitation setting, where the goal is to mimic the behavior as closely as possible given the expert demonstrations; (2) the sparsified reward tasks, where the reward of one is given only when the agent reaches the goal state. We show that the rewards obtained through AILOT relabelling in both settings are descriptive enough to recover the demonstrated policy.
Dataset
OTR, $K=1$ AILOT, $K=1$ OTR, $K=5$ AILOT, $K=5$
@@ 5.3 Results
Table 3: Normalized scores for D4RL locomotion tasks with $K=5$ expert trajectories. The highest scores are highlighted.
Offline Imitation Learning.
We compare AILOT performance on IL tasks in the offline setting. D4RL MuJoco locomotion datasets are used as the main source of offline data with the original reward signal and the action labels discarded, and the expert trajectories chosen for each task as the best episodes achieving maximal return. Results for D4RL locomotion tasks are presented in Table 1. AILOT achieves the best performance in $7$ out of $9$ benchmarks, when compared to OTR and CLUE. Note that, unlike the CLUE method, our algorithm does not rely on the action labels. Empirically, we confirm that the optimal geometrically aware map between the expert and the agent trajectories in the ICVF latent space gives reasonable guidance for the agent to learn properly.
We merge expert demos with unlabeled offline data and train ICVF on the resulting dataset, requiring only $\sim 250$k steps to provide task-agnostic useful spatial representations of states dynamics in the form of intents. Such ICVF representations can be used for finding optimal coupling between expert and agent transitions in the intents space, thus ‘pulling’ the agent towards the demonstrated behavior via intrinsic rewards as described in Section 4. We chose Euclidean distance as a measure of similarity in the intent latent space and perform exponential scaling of rewards according to Eq. (6) in order to maintain them in the appropriate range.
Also, we include additional imitation results when a custom demonstrations is provided, as in Figures 2 and 4, where the Hopper expert makes a backward flip and the HalfCheetah stands upwards respectively. Note that our agent’s dataset initially consists of only the random behaviors. Figures 3 and 5 show how the randomly initialized agent learns to recover the desired behavior through the AILOT relabelling.
Sparse-Reward Offline RL Tasks.
Next, we evaluate AILOT on several sparse D4RL benchmarks (namely, AntMaze-v2 and Adroit-v2). To obtain expert trajectories, we consider only those episodes that accomplish the goal task, dismissing all the others. As stated in the previous Section, ICVF provides structured representations in the latent space, making it easy to distinguish between any two distinct trajectories.
Table 2 compares the performance of AILOT+IQL to OTR+IQL,  CLUE+IQL, when only a single demonstration trajectory is available, and to the original IQL. We observe that AILOT outperforms the current state-of-the art results, with the hardest antmaze-large-diverse task showing the most remarkable margin. This proves the ability of AILOT to employ the expressive representations from ICVF through the optimal transport for functional learning in the sparse tasks as well.
5.4 Ablation Studies
Varying the number of expert trajectories.
We investigate whether performance of learned behavior tends to improve with increased number of provided expert trajectories. Table 3 shows overall performance for varying number of expert trajectories from $K=1$ to $K=5$ across OTR and AILOT. However, we observe that performance across both algorithms improves slightly. We make comparison with OTR here because it’s the most similar to ours (it also uses OT for reward labelling and IQL for RL problem solution). Still, AILOT achieves better normalized scores than OTR, thus proving that alignment in intents representation space improves intrinsic rewards labelling in comparison with similar rewards but with pairwise distances between original states.
Intents distance dependence on the steps count.
In the Method section, we have already mentioned that $\mathbb{E}_{t}\|\psi(s_{t+k})-\psi(s_{t})\|^{2}$ is near-linearly dependent on the steps count between the states ($k$). This important property plays a constructive role in the cost matrix in the OT problem (8), and ultimately, gives a good estimate of the distance between the trajectories. Empirical evidence of the near-linear dependence is presented in Figure 6. Here, we also show that the pairwise distances between the original states $\mathbb{E}_{t}\|s_{t+k}-s_{t}\|^{2}$ have no such feature and, thus, their direct use as a cost function is less preferable.
6 Discussion
In our work, we introduced a novel non-adversarial method for extracting expert’s reward function for offline imitation learning, called AILOT. We empirically show that it surpasses the state-of-the art results both in the sparse-reward RL tasks and in the imitation learning setting. Our method removes the traditional constraint, introduced in the previous offline imitation learning algorithms, where the action labels used to be a necessity. To the contrary, AILOT can mimic the expert behaviour without knowing its action labels, and without the ground truth rewards. Moreover, we show that the intrinsic rewards, distilled by AILOT, could be used to efficiently boost the performance of other offline RL algorithms, thanks to the proper alignment to the expert intentions via the optimal transport.
AILOT assumes there is a sufficient access to a large number of unlabeled trajectories of sub-optimal quality. In our work, we have focused on the expert behaviours, typically considered in the offline RL publications: popular synthetic environments with some comprehensible expert movements and, consequently, some sufficiently ‘intuitive’ intentions.
If, however, the expert has several goals or performs a vague action, the imitation efficiency could, of course, drop (because the expert intents may no longer be transparent to the agent). While such a trait would be on par with the way humans learn a certain skill by observing an adept, weighing the hierarchy of multiple possible intents in AILOT could prove useful to further regularize the learning dynamics in such uncertain scenarios. Notably, such an add-on could be also done in an unsupervised manner, so that the annotations would remain unneeded in the end-to-end implementation.
Another direction of future work is to venture into the cross-domain imitation. Based on the results observed here, it should be possible to generalize AILOT to handle the transition shift between the expert and the agent in the presence of larger mismatches pertinent to the different domains. ICVF is agnostic to the domain setting, making us believe that the cross-domain extension is feasible.
In conclusion, the development of AILOT sets a robust benchmark for future generalizations, enhancing ongoing research in crafting generalist agents with comprehensive knowledge distillation capabilities.

### v2

@@ Appendix B Details about the Baseline Models
  • IQL (Kostrikov et al. (2021)) is state-of-the-art offline RL algorithm, which avoids querying out of the distribution actions by viewing value function as a random variable, where upper bound of uncertanity is controlled through expectile of distribution. In our experiments, evaluation is made using ground-truth reward from D4RL tasks.
  • OTR (Luo et al. (2023)) is a reward function algorithm, where reward signal is based on optimal transport distance between states of expert demonstration and reward unlabeled dataset.
  • CLUE (Liu et al. (2023)) learns VAE calibrated latent space of both expert and agent state-action transitions, where intrinsic rewards can be defined as distance between agent and averaged expert transition representations.
  • IQ-Learn (Garg et al. (2021)) is an imitation learning algorithm, which implicitly encodes into learned inverse Q-function rewards and policy from expert data.
  • Diffusion-QL (Wang et al. (2022)) is an state-of-the-art offline RL algorithm, which models learning policy as conditional diffusion model in order to effectively increase expressiveness and provide more flexible regularization towards behavior policy, which collected dataset.
  • SQIL (Reddy et al. (2019)) proposes to learn soft Q-function by setting expert transitions to one and for non-expert transitions to zero.
  • ORIL (Zolna et al. (2020)) utilizes discriminator network which distinguishes between optimal and suboptimal data in mixed dataset to provide reward relabelling through learned discriminator.
  • SMODICE (Ma et al. (2022a)) offline state occupancy matching algorithm, which solves the problem of IL from observations through state divergence minimization by utilizing dual formulation of value function.
Appendix C Extra ablation study: Varying Number of Expert Trajectories
In this section, we report an ablation study on varying the number of provided expert trajectories, $K$. Table 7 compares AILOT and OTR for $K=1,5$. We observe that the extraction of intent representations is crucial for a good performance. We also observed a negligible performance improvement when the number of provided expert trajectories exceeds certain threshold ($K\geq 10$).
Dataset/Method OTR AILOT OTR AILOT
# of expert trajectories $K=1$ $K=1$ $K=5$ $K=5$
@@ Appendix C Extra ablation study: Varying Number of Expert Trajectories
Table 6: Normalized scores for D4RL locomotion tasks with varying number of expert trajectories. The highest scores are highlighted.
Also, we performed several experiments with modified cost function in OTR method, i.e., exactly the same as in our Eq. 10 with varying levels of $k$, which corresponds to intents shift in Algorithm 1. We observe that the lookahead parameter introduces only negligible improvements to OTR (up to $+0.5$), which is an indicator that the temporal grounding of OT is completely lacking in OTR.
Dataset/Method OTR ($k=2$) AILOT
halfcheetah-medium-v2 $43.3$ $\pm\>0.1$ (+0.1) $47.7$± $\pm\>0.2$
halfcheetah-medium-replay-v2 $41.6$± $\pm\>0.8$ (+0.4) $42.4$± $\pm\>0.2$
halfcheetah-medium-expert-v2 $89.8$ $\pm\>3.0$ (+0.2) $92.4$± $\pm\>1.5$
hopper-medium-v2 $79.3$ $\pm\>5.5$ (+0.6) $82.2$± $\pm\>5.6$
hopper-medium-replay-v2 $85.1$ $\pm\>2.6$ (+0.3) $98.7$± $\pm\>0.4$
hopper-medium-expert-v2 $93.5$ $\pm\>20.6$ (+0.2) $103.4$± $\pm\>5.3$
Table 7: Normalized scores for D4RL locomotion tasks with varying number of expert trajectories. The highest scores are highlighted.
Figure 5: Top: HalfCheetah expert sample trajectory, performing an upwards standing which the agent should imitate; Bottom: HalfCheetah agent successfully performs imitation of standing upwards from the observations via AILOT.
Appendix D Additional Comparison Experiments
We also compare the performance of AILOT to Diffusion-QL and Behavior Cloning (10%), which are shown in Table 8. We evaluated Diffusion-QL with recommended parameters from the original paper and tested performance on sparse-reward AntMaze tasks.
Dataset DiffusionQL AILOT
antmaze-large-diverse $56.6$ $\pm\>7.6$ $66.6$ $\pm\>3.1$
antmaze-large-play $46.4$ $\pm\>8.3$ $57.6$ $\pm\>6.6$
Dataset BC-10 AILOT
halfcheetah-medium-v2 $42.5$ $47.7$ $\pm\>0.2$
halfcheetah-medium-replay-v2 40.6 $42.4$ $\pm\>0.2$
halfcheetah-medium-expert-v2 92.9 $92.4$ $\pm\>1.5$
hopper-medium-v2 56.9 $82.2$ $\pm\>5.6$
hopper-medium-replay-v2 75.9 $98.7$ $\pm\>0.4$
hopper-medium-expert-v2 110.9 $103.4$ $\pm\>5.3$
walker2d-medium-v2 75.0 $78.3$ $\pm\>0.8$
walker2d-medium-replay-v2 62.5 $77.5$ $\pm\>3.1$
walker2d-medium-expert-v2 109.0 $110.2$ $\pm\>1.2$
Table 8: Peformance of AILOT in comparison to Diffusion-QL on sparse-reward AntMaze task and BC-10 results.
Appendix E Ablating $\alpha$ and $\tau$
In the current section we provide expriments on how hyperparameters $\alpha$ and $\tau$ are influencing overall performance. Provided results are similiar to those reported in (Luo et al., 2023).
Dataset AILOT($\alpha=5$, $\tau=0.5$) AILOT($\alpha=1$, $\tau=1$)
halfcheetah-medium-v2 47.7 ± $\pm\>0.35$ 45.6 ± $\pm\>0.3$
halfcheetah-medium-replay-v2 42.4 ± $\pm\>0.8$ 40.2 ± $\pm\>0.5$
halfcheetah-medium-expert-v2 92.4 ± $\pm\>1.5$ 89.4 ± $\pm\>0.7$
walker2d-medium-v2 78.3 ± $\pm\>0.8$ 74.3 ± $\pm\>0.4$
walker2d-medium-replay-v2 77.5 ± $\pm\>3.1$ 71.7 ± $\pm\>2.1$
walker2d-medium-expert-v2 110.2 ± $\pm\>1.2$ 96.7 ± $\pm\>1.0$
Table 9: How the hyperparameters $\alpha$ and $\tau$ influence the performance.
Appendix F Description of Evaluated Environments
Gym-MuJoCo locomotion are the most commonly used tasks for the evaluation of performance of RL algorithms. D4RL provides precollected datasets of different quality, corresponding to training policies evaluated after training for certain number of steps.
Gym-Locomotion.
Locomotion tasks consist of four environments: Hopper, Walker2d, HalfCheetah, and Ant. Provided datasets from D4RL come in "-expert", "-replay", and "-random" splits. Where "-expert" split corresponds to high-reward trajectories, whilst the others contain a mixture of roll-outs from partially trained policies or even some random transitions. All of the datasets are obtained with the SAC algorithm.
Manipulation
In addition to the locomotion experiments, we test the proposed approach on Adroit environments. Adroid involves controlling a 24-DoF Hand robot tasked with opening a door, hammering a nail, picking up and moving a ball. Two types of offline data from D4RL are considered: human demonstrations (with "-human" split), which were obtained from logged experiences of humans and data obtained from a sub-optimal policy (with "-cloned" split), obtained from the imitation policy.